OpenStudy (anonymous):

Eliminate the parameter. x = 5t, y = t + 8

OpenStudy (loser66):

@worne001

OpenStudy (amistre64):

lol, let t= x/5

OpenStudy (amistre64):

x = 5(x/5) = x

OpenStudy (loser66):

@amistre64 the asker is not here to get help, so that's why I didn't start yet. However, yours is not right, hehe

OpenStudy (anonymous):

\[t=y-8\] \[x=5(y-8)\] \[x=5y-40\] \[5y=x+40\] \[y=\frac{ x+40 }{ 5 }\]

OpenStudy (amistre64):

mines never right .. but i do believe its correct ;)

OpenStudy (loser66):

surely x =x is aaaaalways right. 100% agree. hihihi....

OpenStudy (loser66):

but it's not what the asker wants, right?

OpenStudy (amistre64):

they want a y=mx+b setup in the end

OpenStudy (amistre64):

there are many ways to approach it .... erics was fine

OpenStudy (loser66):

yeah, I see.

OpenStudy (anonymous):

yes @amistre64 , thank you. I dont understand the process

OpenStudy (amistre64):

t is the parameter .. so we want to form this into a t-less construction

OpenStudy (amistre64):

we can equate "t" from both equations x = 5t ; t = x/5 y = t+8 ; t = y-8 since t has to equal t at all times ... x/5 = y- 8

OpenStudy (anonymous):

Solve either x or y for t. Now you have that variable in terms of t. Go into the other equation and replace the t with what we solved it for.

OpenStudy (anonymous):

okay, I tried that with a practice prob and it worked when I applied it, ty guys

OpenStudy (amistre64):

good luck :)

OpenStudy (anonymous):

I have another problem that has a square root, x = sqrt of t, y=4t+1

OpenStudy (anonymous):

how would I start this process?

OpenStudy (amistre64):

just square to undo a sqrt

OpenStudy (amistre64):

x^2 = t

OpenStudy (anonymous):

how do you sqrt a letter?

OpenStudy (amistre64):

a letter is just a placeholder for "some number"

OpenStudy (amistre64):

its the number (whatever it may be) that is getting sqrted

OpenStudy (anonymous):

so would this be the right answer? y = 4x2 - 1

OpenStudy (amistre64):

well, +1 ... but yes

OpenStudy (anonymous):

why + 1?

OpenStudy (anonymous):

It was +1 to begin with wasn't it?

OpenStudy (amistre64):

because that what they give us: y = 4t + 1

OpenStudy (amistre64):

t = x^2 doesnt alter anything but the "t"

OpenStudy (anonymous):

oh okay I see..

OpenStudy (anonymous):

I appreciate the help.. im no math wiz haha

OpenStudy (anonymous):

there is one more Im having trouble with..